# On the capability of high redshift kSZ measurement with galaxy surveys

### v1

@@ 3.2 Mock kSZ/CMB maps
To construct maps for the primary CMB and the instrument noise, we utilize their angular power spectrum. First, we establish a same two-dimensional gridding with that of the kSZ map in Fourier space. Each point in this Fourier grid is endowed with a complex amplitude of the square root of the angular power spectrum $\sqrt{C_{l,X}}$, $X$ symbolizing either the CMB or noise, and is complemented by a randomly assigned phase. Transforming the map from Fourier space to real space, a map of temperature fluctuations is achieved. Employing the CAMB software [43], we simulate the primary CMB angular power spectrum, while the noise power spectrum is generated based on the instrumental white noise level $\Delta_{T}$
@@
where $l_{\rm beam}=\sqrt{8\ln 2}/\theta_{\rm FWHM}$ and $\theta_{\ \rm FWHM}$ is the beam size, to all three components; kSZ effect, primary CMB and instrument noise. The CMB instrument noise parameters are set to $\Delta_{T}=11\mu K-{\rm arcmin}$, $\theta_{\rm FWHM}=1.26{\rm arcmin}$ for ACT case and $\Delta_{T}=1\mu K-{\rm arcmin}$, $\theta_{\rm FWHM}=1\ {\rm arcmin}$ for CMB-S4 case. In top panels of figure 1, we illustrate the mock kSZ and the primary CMB plus ACT instrumental noise map, to show the relative scale and amplitude of kSZ effect and its measurement noise, in the redshift interval $2.41.6<z<$ 2.4 as an example. The bottom panel shows the angular power spectra of kSZ effect in three redshift bins, $1.00.6<z<$ 1.0, $1.61.0<z<$ 1.6, $2.41.6<z<$ 2.4, primary CMB and the instrument noise for both ACT and the upcoming CMB-S4 experiment ($C^{\rm Noise}_{l}/W^{2}_{\rm beam}(l)$). Notably, the amplitude of the kSZ effect appears lower than the combined intensity of the CMB and noise, by more than two orders of magnitude. This underscores the challenge of distinguishing the kSZ signal against the background of primary CMB fluctuations and instrumental noise, emphasizing the need for robust techniques to extract and interpret the kSZ signal accurately.
3.3 Mock galaxy, cluster catalogs
We generate the galaxy catalog from the FOF group catalog from the simulation and the ELG HOD. In this work we focus on the redshift range $2.40.6<z<$ 2.4. The redshift range is divided into three redshift bins, $10.6<z<$ 1, $1.61<z<$ 1.6, $2.41.6<z<$ 2.4, of which we use different ELG HODs (the details is shown in appendix B). To generate the galaxy catalog mock, we use the following steps:
@@ 3.3 Mock galaxy, cluster catalogs
For the cluster catalog, we use the halos from the simulation whose mass is above a specific mass limitation. The photometric redshift uncertainty is assigned by $\sigma_{z}=0.01(1+z)$. We use this cluster catalog to determine where to stacking the filtered CMB temperature. This cluster sample is assumed to be identified from photometric galaxies using a group finder such as ref. [36]. The clusters can point out the positions where is abundant of free electron and improve the measurement significance.
@@ 4 Forecast kSZ measurements with PFS survey
Figure 2: The correlation functions of the reconstructed density and velocity fields with the true fields from N-body simulation in three redshift bins. The first row shows the correlation function of the reconstructed density and the matter density from simulation. The blue lines represent the density reconstructed from the spectroscopic galaxy catalog with a fiducial PFS galaxy number density, while the green lines represent the density reconstruction with 20% more galaxies. The second row shows the ratio of the green and blue line, elucidating how an augmentation in galaxy number density can potentially enhance the performance of reconstruction. The third and forth rows shows the same analysis but focus on the reconstructed velocity field. For comparison, the correlation parameter between the reconstructed velocity and the true values of individual clusters is from 0.31 to 0.38, depending on the cluster mass and redshift.
@@
And the correlation parameter of the clusters’ LOS velocity
@@ 4.1 Stacking method
The kSZ stacking method involves the stacking of the filtered CMB temperature signals at the position of clusters weighted by LOS velocity of clusters. The LOS velocity is reconstructed using the method outlined in section 2 and a spectroscopic redshift galaxy catalog, as described in section 3.3. This velocity-weighted stacking helps suppress the tSZ effect, dust contamination, and other foregrounds that do not correlate with the LOS velocities of the clusters.
@@
Additionally, the amplitude of the power spectrum of the primary CMB is larger than that of the kSZ effect by more than three orders of magnitude. In larger scales, the CMB component is more dominant, and at the scale of clusters, it exhibits almost none variation. To effectively isolate the kSZ signal from the dominating CMB fluctuations, an aperture photometry (AP) filter is applied at the location of each cluster
@@
where $\vec{\theta_{i}}$ is the angular position of the cluster $i$. The AP filter
@@
designed to be compensated, integrates over an area to zero. This feature enables the filter to cancel out fluctuations that have a wavelength longer than its size. By varying the size of the AP filter from 0.3 to 5 arcmin, it is possible to delineate the kSZ profile around clusters. The stacking estimator is
@@
where $T_{{\rm AP},i}(\theta_{\rm AP})$ is the CMB temperature at the position of the cluster $i$ after applying the AP filter whose size is $\theta_{\rm AP}$, $v_{{\rm rec},i}$ is the reconstructed LOS velocity of the cluster $i$, $v^{\rm rms}_{\rm rec}$ is the root mean square of the reconstructed LOS velocity. And $r_{v}$ is the correlation coefficient of the reconstructed and true los cluster velocity.
We use the Jackknife method [44] to estimate the precision of measurements. The cluster catalog is divided into $N_{\rm JK}$ subsamples. Then, we measure the observable (eq. 4.9) by removing one subsample from the dataset at a time. The Jackknife covariance is
@@
Figure 3 presents our forecast of the kSZ stacking measurement significance across three redshift bins for both the ACT and CMB-S4 surveys. Notably, the CMB-S4 survey offers a significant improvement, approximately a fivefold enhancement compared to the ACT survey. This improvement arises primarily from the superior sensitivity of CMB-S4. When compared to the ACT case, the stacking profiles from CMB-S4 exhibit a higher S/N for $\theta_{AP}<2$ arcmin. However, the characteristic scale of clusters ($\sim 1$ arcmin, corresponding to $l\sim 2\times 10^{4}$) is dominated by instrumental noise for both ACT and CMB-S4.
@@
Our findings further reveal a notable dependence of the S/N on cluster number density. Specifically, increasing the cluster number density within the range of $10^{-4}$ to $7\times 10^{-3}\ [h/{\rm Mpc}]^{3}$ leads to an approximate twofold increase in the S/N. Interestingly, augmenting galaxy density by 20% yielded negligible changes in measurement significance and the correlation coefficient. This aligns with the findings of ref. [25], suggesting that the reconstruction performance saturates when the galaxy number density exceeds $5\times 10^{-5}[h/{\rm Mpc}]^{3}$.
@@
Furthermore, the correlation parameter between true and reconstructed cluster LOS velocity is between 0.31 to 0.38, and we observe a decreasing trend the correlation parameter with increasing cluster mass (in this case, inferred from photometric cluster redshifts). This relationship is reversed if the cluster redshift is spectroscopic instead of photometric. Notably, the LOS velocity amplitude of clusters exhibits no apparent dependence on cluster mass. We therefore conclude that the decreasing correlation for massive clusters arises from the large velocity disparity around them.
@@
Finally, if the cluster number density could reach $1\times 10^{-3}\ [h/{\rm Mpc}]^{3}$, the S/N of the kSZ stacking method for the three redshift bins would be expected to reach 12, 17, and 18 for the ACT case and 52, 75, and 85 for the CMB-S4 case, respectively. Two primary factors influence the stacking measurements in different redshift bins: survey volume and cluster angular radius. For a fixed cluster number density, a larger survey volume translates to a higher S/N. However, at higher redshifts, the clusters with same mass will have a smaller angular radius, making it more susceptible to instrumental noise.
@@ 4.2 Tomography method
The relatively low S/N observed for the ACT survey in the first redshift bin ($10.6<z<$ 1) suggests limitations in current CMB surveys for measuring the kSZ effect via tomography at redshifts below z = 1. In contrast, the S/N ratio exhibits a significant rise for both surveys in the second and the third redshift bin.
This improvement is attributed to several factors. Firstly, at higher redshifts, the comoving number density of free electrons increases, leading to a stronger kSZ signal. Secondly, larger survey volumes encompass a greater number of clusters, contributing to a higher S/N. Thirdly, for the second redshift bin, while a higher galaxy density may not significantly improve velocity reconstruction, it does contribute to a more accurate reconstruction of the underlying density field, which is crucial for tomography measurements. Fourthly, the non-linear components is reduced with the increasing of the redshift and this would also increase the reconstruction performance.
@@
Above all, both stacking and tomography methods in the PFS survey demonstrate powerful performance in high redshift bins. This indicates a significant potential for the PFS survey in measuring the kSZ effect at high redshifts, marking it as a key merit of this observational program. In addition, while the kSZ stacking and tomography estimators are theoretically equivalent to the bispectrum $\langle\delta\delta v\rangle$ (appendix C), they measure the 1h-term and 2h-term components respectively, like correlation function and power spectrum. Moreover, the kSZ tomography measurement only relying on PFS spectroscopic galaxy catalog possess higher S/N than that of stacking method in the two high redshift bins. This is attributed to the tomography signal capturing contributions from free electrons across various cosmic structures, including halos, filaments, and voids, whereas the stacking approach predominantly measures kSZ effects from massive clusters. Additionally, the scale region where the 1h-term dominates ($l\sim 10^{4}$) is severely suffered by instrumental noise, even within the CMB-S4 survey. Lastly, the kSZ tomography is more sensitive to the galaxy number density comparing to stacking method. The increasing of galaxy number density would not only improve the velocity reconstruction performance, but also the reconstructed density field itself. In addition, we find the improvement of the correlation of the reconstructed and true field show up only when $k>0.03h/{\rm Mpc}$ and slightly increase with the increasing of $k$ (figure. 2). Therefore, the tomography method is benefit from both improvement of small scales and of reconstructed density fields.
@@ 5 Conclusion
To forecast the measurement significance, we use a N-body simulation to generate mock kSZ map, galaxy and group catalogs. For the kSZ mock, we assume the baryon distribution trace that of dark matter. This assumption does not account for the redistribution of baryons driven by baryonic processes. These processes, such as supernovae and AGN feedback, can expel baryonic material from clusters into the surrounding environment. This effect is likely a subdominant contributor to forecasting kSZ measurement significance and will be addressed in future discussions.
@@
To generate the realistic PFS galaxy catalog, we employ realistic ELG HODs from a dark halo catalog. Randomly selecting galaxies ensures the galaxy number density aligns with the predicted value from PFS. Our results underscore the advantages of high redshift kSZ measurement due to its substantial signal amplitude and better velocity reconstruction performance in both methods. In the stacking method, while using the PFS galaxy sample to reconstruct velocity, we use another cluster sample to pin down where to stack kSZ temperature. Interestingly, we observe that the S/N is strongly dependent on the number density of the cluster sample rather than the density of PFS galaxies. The tomography method, which only needs the PFS galaxy sample, can detect kSZ effect with the S/N = 8, 32, 28 for ACT survey and 80, 310, 250 for CMB-S4 survey respectively in the three redshift bins $1.00.6<z<$ 1.0, $1.61.0<z<$ 1.6 and $2.41.6<z<$ 2.4.
@@
While the stacking method excels in measuring the detailed electron profile around massive clusters, tomography captures all free electrons in clusters, filaments, and voids, making it sensitive to the global baryon abundance. These two methods are complementary. Leveraging the PFS spectroscopic redshift survey, the kSZ measurement allows the constraint of baryon abundance extension to $z=2.4$. Simultaneously, the kSZ measurement serves as a valuable tool to constrain baryonic process strength, offering insights into the distribution of baryons. This provides an observable means to calibrate sub-grid parameters in hydrodynamic simulations and address one of the most significant theoretical systematic uncertainties in weak lensing measurements.
@@
Furthermore, to enhance the measurement significance beyond redshift $z=1$, the incorporation of other high redshift data such as, the Lyman-$\alpha$ forest and 21cm internsity mapping, could significantly improve the reconstruction performance of peculiar velocities and redshift range. The observed increase in quasar number density facilitates more accessible and accurate density measurements through their absorption lines. But the sampling across the sky can be sparse. This characteristic poses challenges for velocity reconstruction, which we defer to future investigations.

### v2

@@ 3.2 Mock kSZ/CMB maps
To construct maps for the primary CMB and the instrument noise, we utilize their angular power spectrum. First, we establish the same two-dimensional gridding with that of the kSZ map in Fourier space. Each point in this Fourier grid is endowed with a complex amplitude of the square root of the angular power spectrum $\sqrt{C_{l,X}}$, $X$ symbolizing either the CMB or noise, and is complemented by a randomly assigned phase. Transforming the map from Fourier space to real space, a map of temperature fluctuations is achieved. Employing the CAMB software [49], we simulate the primary CMB angular power spectrum, while the noise power spectrum is generated based on the instrumental white noise level $\Delta_{T}$
@@
where $l_{\rm beam}=\sqrt{8\ln 2}/\theta_{\rm FWHM}$ and $\theta_{\ \rm FWHM}$ is the beam size, to the kSZ and primary CMB components. The CMB instrument noise parameters are set to $\Delta_{T}=11\mu K-{\rm arcmin}$, $\theta_{\rm FWHM}=1.26{\rm arcmin}$ for ACT case and $\Delta_{T}=1\mu K-{\rm arcmin}$, $\theta_{\rm FWHM}=1\ {\rm arcmin}$ for CMB-S4 case. In top panels of figure 1, we illustrate the mock kSZ and the primary CMB plus ACT instrumental noise map, to show the relative scale and amplitude of kSZ effect and its measurement noise, in the redshift interval $2.41.6<z<$ 2.4 as an example. The bottom panel shows the angular power spectra of kSZ effect in three redshift bins, $1.00.6<z<$ 1.0, $1.61.0<z<$ 1.6, $2.41.6<z<$ 2.4, primary CMB and the instrument noise for both ACT and the upcoming CMB-S4 experiment ($C^{\rm Noise}_{l}/W^{2}_{\rm beam}(l)$). Notably, the amplitude of the kSZ effect appears lower than the combined intensity of the CMB and noise, by more than two orders of magnitude. This underscores the challenge of distinguishing the kSZ signal against the background of primary CMB fluctuations and instrumental noise, emphasizing the need for robust techniques to extract and interpret the kSZ signal accurately.
3.3 Mock galaxy, galaxy group catalogs
We generate the galaxy catalog from the FOF group catalog from the simulation and the ELG HOD. In this work we focus on the redshift range $2.40.6<z<$ 2.4. The redshift range is divided into three redshift bins, $10.6<z<$ 1, $1.61<z<$ 1.6, $2.41.6<z<$ 2.4, of which we use different ELG HODs (the details are shown in appendix B). To generate the galaxy catalog mock, we use the following steps:
@@ 3.3 Mock galaxy, galaxy group catalogs
For the galaxy group catalog, we use the halos from the simulation whose mass is above a specific mass limitation. The photometric redshift uncertainty is assigned by $\sigma_{z}=0.01(1+z)$. We use this galaxy group catalog to determine where to stacking the filtered CMB temperature. This galaxy group sample is assumed to be identified from photometric galaxies using a group finder such as ref. [41]. The galaxy groups can point out the positions where is abundant of free electron and improve the measurement significance.
@@ 4 Forecast kSZ measurements with PFS survey
Figure 2: The correlation functions of the reconstructed density and velocity fields with the true fields from N-body simulation in three redshift bins. The first row shows the correlation function of the reconstructed density and the matter density from simulation. The blue lines represent the density reconstructed from the spectroscopic galaxy catalog with a fiducial PFS galaxy number density, while the green lines represent the density reconstruction with 20% more galaxies. The second row shows the ratio of the green and blue line, elucidating how an augmentation in galaxy number density can potentially enhance the performance of reconstruction. The third and forth rows shows the same analysis but focus on the reconstructed velocity field. For comparison, the correlation parameter between the reconstructed velocity and the true values of individual galaxy groups is from 0.31 to 0.38, depending on the galaxy group mass and redshift.
@@
And the correlation parameter of the galaxy groups’ LOS velocity
@@ 4.1 Stacking method
Figure 3: The kSZ profile from the stacking measurement (eq. 4.9). The galaxy group mass range is $12.3912.29<{\rm lgM}<$ 12.39 and the redshift range is $2.41.6<z<$ 2.4. The blue points show the case with only kSZ component and the orange points show the case with primary CMB, kSZ and ACT instrument noise components.
@@
The kSZ stacking method involves the stacking of the filtered CMB temperature signals at the position of galaxy groups weighted by their LOS velocity. The LOS velocity is reconstructed using the method outlined in section 2 and a spectroscopic redshift galaxy catalog, as described in section 3.3. This velocity-weighted stacking helps suppress the tSZ effect, dust contamination, and other foregrounds that do not correlate with the LOS velocities of the galaxy groups.
@@
Additionally, the amplitude of the power spectrum of the primary CMB is larger than that of the kSZ effect by more than three orders of magnitude. In larger scales, the CMB component is more dominant, and at the scale of galaxy groups, it exhibits almost none variation. To effectively isolate the kSZ signal from the dominating CMB fluctuations, an aperture photometry (AP) filter is applied at the location of each galaxy group
@@
where $\vec{\theta_{i}}$ is the angular position of the galaxy group $i$. The AP filter
@@
designed to be compensated, integrates over an area to zero. This feature enables the filter to cancel out fluctuations that have a wavelength longer than its size. Note if the size of the AP filter is smaller or comparable to the size of the galaxy groups, the AP filter would subtract the kSZ signal in the annulus region, thereby suppressing the total kSZ signal [50, 51]. By varying the size of the AP filter from 0.3 to 5 arcmin, it is possible to delineate the kSZ profile around galaxy groups. The stacking estimator is
@@
where $T_{{\rm AP},i}(\theta_{\rm AP})$ is the CMB temperature at the position of the galaxy group $i$ after applying the AP filter whose size is $\theta_{\rm AP}$, $v_{{\rm rec},i}$ is the reconstructed LOS velocity of the galaxy group $i$, $v^{\rm rms}_{\rm rec}$ is the root mean square of the reconstructed LOS velocity. And $r_{v}$ is the correlation coefficient of the reconstructed and true los galaxy group velocity.
We use the Jackknife method [52] to estimate the precision of measurements. The galaxy group catalog is divided into $N_{\rm JK}$ subsamples. Then, we measure the observable (eq. 4.9) by removing one subsample from the dataset at a time. The Jackknife covariance is
@@
Figure 3 show the kSZ profile measurement of a galaxy group sample whose mass range is $12.3912.29<{\rm lgM}<$ 12.39 in the third redshift bin with the ACT survey as an example. The kSZ-only case is consistent with the one adding primary CMB and ACT instrument noise within the errorbars, showing the stacking estimate (eq. 4.9) is unbiased. Figure 4 presents our forecast of the kSZ stacking measurement significance across three redshift bins for both the ACT and CMB-S4 surveys. Notably, the CMB-S4 survey offers a significant improvement, approximately a fivefold enhancement compared to the ACT survey. This improvement arises primarily from the superiority sensitivity of CMB-S4. When compared to the ACT case, the stacking profiles from CMB-S4 exhibit a higher S/N for $\theta_{AP}<2$ arcmin. However, the characteristic scale of galaxy groups ($\sim 1$ arcmin, corresponding to $l\sim 2\times 10^{4}$) is dominated by instrumental noise for both ACT and CMB-S4.
$\bar{n}_{g}\times 10^{-4}[{\rm Mpc}/h]^{-3}$
Redshift bin 1 2 3 5 7 9 30 70
$1.00.6<z<$ 1.0 13.30 13.08 12.94 12.75 12.62 12.52 12.02 11.65
$1.61.0<z<$ 1.6 13.13 12.92 12.80 12.62 12.50 12.41 11.95 11.60
$2.41.6<z<$ 2.4 12.84 12.67 12.54 12.39 12.29 12.21 11.80 11.49
Table 2: The minimum masses of galaxy group samples in the three redshift bins. Each column corresponds to a galaxy group number density, as labeled in the second line.
@@
Our findings further reveal a notable dependence of the S/N on galaxy group number density. Specifically, increasing the galaxy group number density within the range of $10^{-4}$ to $7\times 10^{-3}\ [h/{\rm Mpc}]^{3}$ leads to an approximately twofold increase in the S/N. The corresponding minimum masses of these galaxy group samples are shown in table 2. Interestingly, augmenting galaxy density by 20% yielded negligible changes in measurement significance and the correlation coefficient. This aligns with the findings of ref. [27], suggesting that the reconstruction performance saturates when the galaxy number density exceeds $5\times 10^{-5}[h/{\rm Mpc}]^{3}$.
@@
Furthermore, the correlation parameter between true and reconstructed galaxy group LOS velocity is between 0.31 to 0.38, and we observe a decreasing trend in the correlation parameter with increasing galaxy group mass (in this case, inferred from photometric galaxy group redshifts). This relationship is reversed if the galaxy group redshift is spectroscopic instead of photometric. Notably, the LOS velocity amplitude of galaxy groups exhibits no apparent dependence on galaxy group mass. We therefore conclude that the decreasing correlation for massive galaxy groups arises from the large velocity disparity around them.
@@
In addition, if the galaxy group number density could reach $1\times 10^{-3}\ [h/{\rm Mpc}]^{3}$, the S/N of the kSZ stacking method for the three redshift bins would be expected to reach 12, 17, and 18 for the ACT case and 52, 75, and 85 for the CMB-S4 case, respectively. Two primary factors influence the stacking measurements in different redshift bins: survey volume and galaxy group angular radius. For a fixed galaxy group number density, a larger survey volume translates to a higher S/N. However, at higher redshifts, the galaxy groups with same mass will have a smaller angular radius, making it more susceptible to instrumental noise.
Finally, in this prediction analysis, we use an external photometric galaxy group catalog as the position proxies for stacking kSZ effect. However, the certain galaxies, such as LRGs, can serves as proxies of galaxy groups. Thus, the full stacking measurement could just use the PFS LRG catalog. We leave this to future work.
@@ 4.2 Tomography method
The relatively low S/N observed for the ACT survey in the first redshift bin ($10.6<z<$ 1) suggests limitations in current CMB surveys for measuring the kSZ effect via tomography at redshifts below z = 1. In contrast, the S/N ratio exhibits a significant rise for both surveys in the second and the third redshift bins.
This improvement is attributed to several factors. Firstly, at higher redshifts, the comoving number density of free electrons increases, leading to a stronger kSZ signal. Secondly, larger survey volumes encompass a greater number of galaxy groups, contributing to a higher S/N. Thirdly, for the second redshift bin, while a higher galaxy density may not significantly improve velocity reconstruction, it does contribute to a more accurate reconstruction of the underlying density field, which is crucial for tomography measurements. Fourthly, the non-linear components are reduced with the increasing of the redshift and this would also increase the reconstruction performance.
@@
Above all, both stacking and tomography methods in the PFS survey demonstrate powerful performance in high redshift bins. This indicates a significant potential for the PFS survey in measuring the kSZ effect at high redshifts, marking it as a key merit of this observational program. While the kSZ stacking and tomography estimators are theoretically equivalent to the bispectrum $\langle\delta\delta v\rangle$ (appendix C), they measure the 1h-term and 2h-term components respectively, like correlation function and power spectrum. Moreover, the kSZ tomography measurement only relying on PFS spectroscopic galaxy catalog possesses higher S/N than that of stacking method in the two high redshift bins. This is attributed to the tomography signal capturing contributions from free electrons across various cosmic structures, including halos, filaments, and voids, whereas the stacking approach predominantly measures kSZ effects from massive clusters. Additionally, the scale region where the 1h-term dominates ($l\sim 10^{4}$) is severely suffered by instrumental noise, even within the CMB-S4 survey. In addition, the kSZ tomography is more sensitive to the galaxy number density compared to stacking method. The increasing of galaxy number density would not only improve the velocity reconstruction performance, but also the reconstructed density field itself. We find the improvement of the correlation of the reconstructed and true field shows up only when $k>0.03h/{\rm Mpc}$ and slightly increases with the increasing of $k$ (figure. 2). Therefore, the tomography method benefits from both improvement of small scales and reconstructed density fields. Lastly, the instrument noise of CMB-S4 is about ten times smaller than that of ACT. Therefore, for both stacking and tomography methods, the S/N for CMB-S4 is significantly higher than that of ACT.
@@ 5 Conclusion
To forecast the measurement significance, we use an N-body simulation to generate mock kSZ map, galaxy and group catalogs. For the kSZ mock, we assume the baryon distribution trace that of dark matter. This assumption does not account for the redistribution of baryons driven by baryonic processes. These processes, such as supernovae and AGN feedback, can expel baryonic material from galaxy groups into the surrounding environment. This effect is likely a subdominant contributor to forecasting kSZ measurement significance and will be addressed in future work. Here, we just discuss how the baryon feedback would influence our results. The stacking method is more sensitive to the feedback especially when $\theta_{\rm AP}$ is small, because the signal of kSZ effect in groups is proportional to the amount of baryon. Therefore, the S/N is decreasing with a scaling of $f_{\rm gas}$. On the other hand, the stacking measurement and its S/N can provide a constraint on $f_{\rm gas}$ and then the baryon feedback model.
@@
To generate the realistic PFS galaxy catalog, we employ realistic ELG HODs from a dark halo catalog. Randomly selecting galaxies ensures the galaxy number density aligns with the predicted value from PFS. Our results underscore the advantages of high redshift kSZ measurement due to its substantial signal amplitude and better velocity reconstruction performance in both methods. In the stacking method, while using the PFS galaxy sample to reconstruct velocity, we use another galaxy group sample to pin down where to stack kSZ temperature. Interestingly, we observe that the S/N is strongly dependent on the number density of the galaxy group sample rather than the density of PFS galaxies. The tomography method, which only needs the PFS galaxy sample, can detect kSZ effect with the S/N = 8, 32, 28 for ACT survey and 80, 310, 250 for CMB-S4 survey respectively in the three redshift bins $1.00.6<z<$ 1.0, $1.61.0<z<$ 1.6 and $2.41.6<z<$ 2.4.
@@
While the stacking method excels in measuring the detailed electron profile around massive galaxy groups, tomography captures all free electrons in groups, filaments, and voids, making it sensitive to the global baryon abundance. These two methods are complementary. Leveraging the PFS spectroscopic redshift survey, the kSZ measurement allows the constraint of baryon abundance extension to $z=2.4$. Simultaneously, the kSZ measurement serves as a valuable tool to constrain baryonic process strength, offering insights into the distribution of baryons. This provides an observable means to calibrate sub-grid parameters in hydrodynamic simulations and address one of the most significant theoretical systematic uncertainties in weak lensing measurements.
@@
Furthermore, to enhance the measurement significance beyond redshift $z=1$, the incorporation of other high redshift data such as, the Lyman-$\alpha$ forest and 21cm intensity mapping, could significantly improve the reconstruction performance of peculiar velocities and redshift range. The observed increase in quasar number density facilitates more accessible and accurate density measurements through their absorption lines. But the sampling across the sky can be sparse. This characteristic poses challenges for velocity reconstruction, which we defer to future investigations.